OpenStudy (anonymous):

Find the difference of (−4x2 − 7x3 − 2) − (8x3 + 3x − 7x2). Be sure to show all of your work in a neat and organized fashion. Use the ^ (shift 6) symbol to show an exponent. For example, x squared can be written as x^2.

OpenStudy (anonymous):

So what is it about this problem that you need help with?

jimthompson5910 (jim_thompson5910):

(-4x^2 - 7x^3 - 2) - (8x^3 + 3x - 7x^2) -4x^2 - 7x^3 - 2 - 8x^3 - 3x + 7x^2 ... Distribute Now combine like terms.

OpenStudy (anonymous):

okay so -7x^3-5x^2-3x-2 ?

OpenStudy (anonymous):

Is that correct? :/

jimthompson5910 (jim_thompson5910):

- 7x^3 - 8x^3 = ???

OpenStudy (anonymous):

-15x^3

jimthompson5910 (jim_thompson5910):

-4x^2 + 7x^2 = ???

OpenStudy (anonymous):

\[\left(-4 x^2-7 x^3-2\right)-\left(8 x^3+3 x-7 x^2\right)=-15 x^3+3 x^2-3 x-2 \]

OpenStudy (anonymous):

-15x^3